OpenStudy (cherrilyn):

Write in Summation Notation:

OpenStudy (cherrilyn):

1/9 + 1/16 + 1/25 + 1/36 + ......... I know it has something to do with a square in the denominator..

OpenStudy (anonymous):

1/n^2 , from 3-infinity

OpenStudy (anonymous):

\[\sum_{1}^{\infty}1/(n+2)^{2}\]

OpenStudy (anonymous):

\[\sum_{n=3}^{\infty} \frac{1}{n^2} \]

OpenStudy (anonymous):

yeah

OpenStudy (cherrilyn):

which one?

OpenStudy (anonymous):

both are same

OpenStudy (cherrilyn):

oh okay thank you :)

OpenStudy (cherrilyn):

what about for 1 - 1/3 + 1/5 - 1/7 +.......

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (-1)^n \frac{1}{2n+1}\]

OpenStudy (cherrilyn):

and 125/9 + 625/16 + 3,125/25 + 15,625/36 . I know the bottom is n^2 but what's the top?

OpenStudy (anonymous):

the top are factgorials im fairly sure

OpenStudy (anonymous):

5!, 6! etc

OpenStudy (anonymous):

\[\sum_{n=3}^{ \infty} \frac{ (n+2)!}{n^2} \]

OpenStudy (anonymous):

wait, thats not right :|

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its multiples of 5 on the top

OpenStudy (anonymous):

its powers of 5

OpenStudy (anonymous):

5^n / n^2

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

n=3 to infinity

OpenStudy (anonymous):

\[\sum_{n=3}^{\infty} 5^n \frac{5^2}{n^2} \]

OpenStudy (anonymous):

\[\sum_{n=3}^{\infty} \frac{5^{n+2}}{n^2}\]

OpenStudy (anonymous):

ur wrong mate

OpenStudy (anonymous):

if it starts from 3 then its 5^n

OpenStudy (anonymous):

no, thats not it :|

OpenStudy (anonymous):

yeh whatever

OpenStudy (cherrilyn):

?

OpenStudy (anonymous):

\[\sum_{n=3}^{\infty}5^{n}/n ^{2}\]

OpenStudy (anonymous):

thats ur answer

OpenStudy (cherrilyn):

you're right ! Thanks :D Saved me

OpenStudy (anonymous):

no prob...

OpenStudy (anonymous):

hey could you look at myt problem pleaseeee